# A new indeterminacy-based quantum theory

### v4

@@ Abstract
I propose a novel (interpretation of) quantum theory, which I will call Environmental Determinacy-based or EnD Quantum Theory (EnDQT). In contrast to the well-known quantum theories, EnDQT has the benefit of not adding hidden variables, and it is not in tension with relativistic causality by providing a local causal explanation of quantum correlations without measurement outcomes varying according to, for example, systems or worlds. It is conservative, and so unlike theories such as spontaneous collapse theories, no modifications of the fundamental equations of quantum theory are required to establish when determinate values arise, and in principle, arbitrary systems can be in a superposition for an arbitrary amount of time. According to EnDQT, some systems acquire determinate values at some point, and the capacity to have and give rise to determinate values through interactions propagates to other systems via local interactions. When systems are isolated from the systems that belong to these chains of interactions, they can, in principle, evolve unitarily indefinitely. EnDQT provides a series of novel empirical posits that may distinguish it from other interpretations of quantum theory. Furthermore, via its characterization of the systems that start these chains of interactions, it provides possible payoffs to other areas of physics and their foundations, such as cosmology, thermodynamics, and statistical mechanics.
1 Introduction
@@
The measurement problem [^1] See, e.g., Maudlin (1995) and Myrvold (2018). arises from interactions in quantum theory (QT), which, without introducing some extra assumptions, can lead the quantum state of a macroscopic system to be in a superposition, where the latter doesn’t correspond to a physical magnitude with determinate values. However, we know from classical physics and experimental evidence that this can’t be the case at macroscopic scales.
In the search for a solution to this problem, as far as we know, a conservative approach should seek to fulfill the following desideratum:
@@
*) A precise criterion for when determinate values arise that doesn’t modify the fundamental equations of QT like spontaneous collapse theories or postulates a special force that causes such collapse like gravitational collapse theories. [^2] See, e.g., Ghirardi & Bassi (2020), Oppenheim (2023), and Diósi (1995). This is because we currently have no clear evidence for that.
Similarly, given some adequate conditions, it’s plausible to consider that any system could, in principle, evolve unitarily indefinitely regarding any physical degree of freedom. So, this conservative solution should also fulfill the desideratum of
@@
**) Allowing for any system to, in principle, be in a superposition of quantum states associated with any physical magnitude for an arbitrary amount of time.
@@
I will consider that an approach that fulfills *) and **), fulfills UT). A conservative strategy to fulfill UT) should aim to allow, in some circumstances, for any system to be placed in a superposition that corresponds to any physical magnitude for an arbitrary amount of time, even large target systems interacting with equally large systems. Then, perhaps consider that only interactions with specific systems lead a target system $S$ to have determinate values, where these interactions are described quantum mechanically via decoherence. [^3] See, e.g., Schlosshauer (2007) and references therein. Decoherence doesn’t modify the fundamental equations of QT as stated by Dirac and von Neumann [^4] See, e.g., Barrett (2019, Ch. 4). and thus allows for a conservative approach. Furthermore, we have evidence that interactions between quantum systems involving decoherence play some role in giving rise to determinate outcomes. But interactions with which specific systems? Again, we could be conservative and appeal to interactions only with systems that were decohered shortly before the decohering interaction with $S$ started, where these later systems were decohered by other systems that were also previously decohered, and so on. However, this idea seems to start giving rise to an infinite regress and some vagueness regarding the details of these interactions.
@@
To deal with these issues, we could appeal to some plausible special systems that establish when these interactions began, a more precise structure that represents such interactions, and simple and conservative rules that establish how determinate values arise from them. Environmental Determinacy-based Quantum Theory (EnDQT) will pursue this strategy by introducing a network structure whose edges represent interactions between certain systems represented via decoherence, and that roughly establishes when these interactions give rise to them having determinate values. These interactions form what I will call stable determination chains (SDCs). Furthermore, SDCs started somewhere. As I will argue, the first systems with determinate values arose in the past through special systems. These systems started chains of local interactions over time and space, which are the SDCs. By interacting with these systems, a system acquires a determinate value of an observable during these interactions and the capacity to give rise to other systems having determinate values in interactions with them, which allows these later systems to lead other systems to have determinate values, and so on. So, these chains allow determinate values to propagate between systems and persist over spacetime, where it is indeterministic which value will arise among the possible ones. These interactions are modeled via decoherence; thus, as I have said, they don’t lead to any modification of fundamental equations of QT. The systems that don’t belong to this network or don’t interact with it at some point can, in principle, unitarily evolve indefinitely.
I will argue for one possibility for what these systems that start SDCs could be via inflation, which is arguably the dominant paradigm in modern cosmology. This would provide a new role for the systems that start inflation. I will also argue that this assumption regarding these systems has diverse philosophical advantages and possible payoffs to other areas of physics, such as grounding the state of the beginning of the universe, the source of temporal asymmetries, the role of the inflaton field, and its effects in something more fundamental.
Furthermore, another important desideratum that a conservative approach to QT should achieve is
LC) In the domains where we know where to apply QT, not being in tension with relativity by not favoring a reference frame or leading to action at a distance like, for example, Bohmian mechanics, [^5] See, e.g., Goldstein (2021). and don’t add hidden variables that lead to retrocausality or superdeterminism. [^6] See, e.g., Hossenfelder & Palmer (2020) and Friederich & Evans (2019).
For EnDQT to achieve LC), first, I will argue that it is able to deal with Bell’s theorem by providing a local explanation of quantum correlations via Quantum Causal Models (section 3). [^7] See, e.g., Costa & Shrapnel (2016), Allen et al. (2017), and Barrett et al. (2019).
@@
Second, EnDQT will adopt a perspective on quantum states where they don’t literally and directly represent some physical entity; instead, together with other elements of the theory, such as observables and networks representing SDCs, they help make inferences, gain knowledge about and indirectly represent how systems evolve and affect each other, how SDCs evolve, when systems acquire or not determinate values, how systems evolve outside interactions, etc.
@@
So, contrary to spontaneous collapse theories, which reify the quantum state, there is no literal physical collapse of quantum states in a superposition, which can be highly entangled and non-local. Instead, there is an epistemic local state update of the original state of the target system that can be implemented upon decoherence of this system by its environmental systems that belong to SDCs, under their local interactions. These interactions give rise indeterministically to these systems having a determinate value. So, given that EnDQT doesn’t reify the quantum state, in the Bell-type scenarios, the measurement of Alice doesn’t non-locally affect the space-like separated Bob and vice versa.
@@
This leads to the view that considers that decoherence shouldn’t be interpreted as representing a process of branching of the wave-function/quantum states but rather as a process in which, under local interactions, an environmental system that belongs to an SDC gives rise to another system having determinate values indeterministically. Furthermore, MWI-like views consider that decoherence in large enough regions of spacetime establishes criteria for systems to have determinate values of an observable. However, as we will see in the next section, EnDQT considers that such criteria form necessary but not sufficient conditions for determinate values to arise. This is because it matters if the environmental systems involved in this process belong to an SDC.
A common way to fulfill UT) and LC) is by adopting a relationalist interpretation in which the outcomes of Alice or Bob are relative to, for example, worlds, private perspectives, environments, simultaneity hyperplanes, etc. [^8] See, e.g., Wallace (2012), Di Biagio & Rovelli (2021), Healey (2017), Dieks (2019), and Ormrod & Barrett (2024). However, for EnDQT these outcomes and all physical states will be absolute/non-relationalist. So, EnDQT,
NR) Doesn’t adopt a relationalist interpretation of QT.
@@
Given the well-known issue of probabilities of the MWI, [^9] See, e.g., Albert (2010) and Price (2010). this might be deemed a desideratum and advantage of this view. Also, it’s unclear if relationalism in a single-world is desirable. Furthermore, as I will argue (section 3), to my knowledge, EnDQT is currently the only QT that doesn’t modify the fundamental equations of standard QT, [^10] It might be objected that quantum causal models modify the fundamental equation of QT. Note that I don’t regard these models as a modification of equations of standard QT, but as a generalization or variation. that is able to provide a local, non-hidden variable, non-relationalist common cause explanation of quantum correlations like the ones in Bell scenarios and the so-called extended Wigner’s friend scenarios. [^11] I will focus on the scenarios from Bong et al. (2020) and Brukner (2018). [^12] Contrary to the suggestions by others (e.g., Cavalcanti & Wiseman, 2021, Schmid et al., 2023, and Yīng et al., 2023), we don’t need to adapt QCMs to a relationalist approach, which might be considered as an advantage of this view.
So, EnDQT should be considered a conservative approach to QT because, as I have said, it doesn’t modify the fundamental equations of standard QT, which I think is a virtue. Furthermore, it has the great benefit of fulfilling all of the above desiderata and providing other payoffs. I will start by explaining the basics of EnDQT, argue that it provides UT) and start building the case that it provides LC) and NR) (section 2). In section 3, I will argue that EnDQT provides LC) and NR) by showing that it provides a non-relational, local, and non-superdeterministic/non-retrocausal explanation of quantum/Bell-type correlations. In section 4 and throughout the text, I will suggest future developments. To simplify, I will mostly assume non-relativistic QT and the Schrödinger picture Hilbert space-based finite dimensional QT.
@@ 2 Main features: four conditions and two hypotheses
I will start by presenting the main features of this view and show why it fulfills UT). Also, I will begin building the argument for why it allows for LC) and NR). The main features of EnDQT presented here [^13] We will see that there are other versions of EnDQT if we vary these hypotheses and conditions, but we will see that this version so far is enough to fulfill the desiderata of Section 1. involve four conditions and two hypotheses about how the capacity to have determinate values, and to provide other systems that capacity, spreads through interactions. I will start by explaining the four conditions, which will require that I give an account of the role of quantum states (which was already given in the previous section), systems, interactions, decoherence, and how having a determinate value and allowing other systems to have determinate values propagate via interactions inferred via decoherence, where specific chains of interactions are formed. Afterward, I will present the two hypotheses that support EnDQT, where the second hypothesis involves an account of what the systems that start SDCs could be. I will mention some empirical predictions that EnDQT provides throughout this section.

### v5

@@ Abstract
I propose a novel interpretation of quantum theory, which I will call Environmental Determinacy-based (EnDQT). In contrast to the well-known interpretations of quantum theory, EnDQT has the benefit of not adding non-local, superdeterministic, or retrocausal hidden variables. Also, it is not in tension with relativistic causality by providing a local causal explanation of quantum correlations. Furthermore, measurement outcomes don’t vary according to, for example, systems or worlds. It is a conservative QT in the sense that, unlike theories such as spontaneous collapse theories, no modifications of the fundamental equations of quantum theory are required to establish when determinate values arise. Moreover, in principle, arbitrary systems can be in a coherent superposition for an arbitrary amount of time. According to EnDQT, at a certain stage of the evolution of the universe, some systems acquire the capacity to have and give rise to other systems having determinate values through an indeterministic process. Furthermore, this capacity propagates via local interactions between systems. When systems are isolated from others that have this capacity, they can, in principle, evolve unitarily indefinitely. EnDQT may provide payoffs to other areas of physics and their foundations, such as cosmology, via the features of the systems that start the chains of interactions.
1 The measurement problem and EnDQT’s desiderata
@@
The measurement problem [^1] See, e.g., [64, 71]. can be seen as arising from interactions in quantum theory (QT), which, without introducing additional assumptions, lead to the quantum state of a macroscopic system being in a superposition. This superposition does not correspond to a physical magnitude with determinate values. However, we know from classical physics and experimental evidence that this cannot be the case at macroscopic scales.
One way to see this problem more concretely is by first assuming the so-called Eigenstate-Eigenvalue Link (EEL). The EEL gives a clear way of making sense of how the mathematical framework of QT represents physical magnitudes and it played an important role in interpreting QT, especially within the more orthodox interpretations. [^2] See [42] for a historical overview of the importance of this link. According to this link:
A system S has a determinate value q of an observable O if and only if the quantum state of S is in eigenstate of $O$ with an eigenvalue $q$.
This link often leads to the assumption that if the quantum state of $S$ is not an eigenstate of some observable, $S$ has, in a sense, an indeterminate value of that observable.
Now, let’s consider a quantum system $S$ and a measurement device $M$, where $M$ will measure a certain property of $S$. The system $S$ is initially in a superposition of quantum states:
$|\psi\rangle_{S}=\sum_{i}c_{i}|\psi_{i}\rangle_{S}$ (1)
where $\{|\psi_{i}\rangle_{S}\}$ are the eigenstates of the observable $\hat{O}$ of the system, with corresponding eigenvalues $\{o_{i}\}$. Furthermore, the measurement device $M$ is initially in a ready state $|m_{0}\rangle_{M}$. States $\{|m_{i}\rangle_{M}\}$ of the device $M$ are eigenstates of the observable $\hat{M}$ that represents the different measurement records of $M$ when it interacts with $S$.
During the measurement process, the system $S$ interacts with the device $M$. According to the standard rules for interactions in unitary QT, this interaction results in the quantum state of the measurement device becoming correlated with the eigenstates of the observable $\hat{O}$, leading to the following entangled state with $c_{i}\neq 0$:
$|\Psi\rangle_{SM}=\sum_{i}c_{i}|\psi_{i}\rangle_{S}|m_{i}\rangle_{M}.$ (2)
Given the EEL, this superposition implies that the system $S$ doesn’t have a determinate value of observable $\hat{O}$, as well as the observable $\hat{M}$. Now, the problem is that this is contrary to what our measurement records indicate, as well as to what classical physics successfully predicts at these scales. We might choose to deal with this issue by assuming the so-called collapse postulate:
When the system S is measured, the state of S transitions stochastically to one the eigenstates of $\hat{O}$, where the probability for such transition is given by the Born rule.
However, it is unclear precisely when to apply this postulate and how to ground it because it is unclear what constitutes a measurement. This issue, in a nutshell, forms the measurement problem.
Diverse quantum theories (QTs) were developed to deal with the measurement problem. Standard QT, whose basic axioms were initially developed by Dirac and von Neumann,333[35, 93]. See Appendix A, and e.g., Chapter 4 in [6] for an introduction., has enjoyed incredible empirical success. Given the success of this theory, a plausible desideratum for a satisfactory quantum theory that solves the measurement problem is to be as conservative as possible, staying as close as possible to standard QT. In the search for a solution to this problem, given the empirical success of QT, I think we should adopt a conservative approach. This conservative approach should seek to fulfill the following desideratum:
@@
*) A precise criterion for when determinate values arise that doesn’t modify the fundamental equations of QT like spontaneous collapse theories or postulates a special force that causes such collapse like gravitational collapse theories. [^4] See, e.g., [41, 74, 34].
This desideratum is conservative because we currently have no clear evidence that we need to modify the equations of QT or posit that a force like gravity causes the collapse. Similarly, current evidence doesn’t point towards limits regarding the scales in which we can place systems in a coherent superposition so that they give rise to interference phenomena if they are sufficiently isolated from their environment. In other words, it is plausible to consider that any system or systems in a superposition of quantum states could, in principle, evolve unitarily indefinitely regarding any dynamical physical magnitude. So, this conservative solution should also fulfill the desideratum of
@@
**) Allowing for any system to, in principle, be in a coherent superposition of quantum states associated with at least any dynamical physical magnitude for an arbitrary amount of time.
I will consider that an approach that fulfills *) and **) fulfills what I will call the desideratum UT) for unitarity.
@@
Decoherence theory suggests that interactions between a $target$ system $S$ with a certain kind of system $E$, which includes measurement devices, lead $S$ to have determinate values. [^5] See, e.g., [89] and references therein. Furthermore, we have evidence that interactions between quantum systems involving decoherence play some role in giving rise to determinate outcomes. [^6] Ibid. Decoherence doesn’t modify the fundamental equations of standard QT and thus allows for a conservative approach that fulfills *). Therefore, a conservative strategy fulfilling *) should consider that only interactions with systems $E$ lead a target system $S$ to have determinate values, where these interactions are described quantum mechanically via decoherence.
But interactions with which systems $E$, which kinds of interactions? We don’t want just to vaguely claim that these would just be interactions with systems that constitute the “large environment” of a system $S$ that undergoes decoherence (as it is traditionally claimed). First, it is unclear which systems we are talking precisely about since the environment of a system can involve many systems. Second, considering the desideratum **), we aim to ensure that any systems, including those interacting with their environment, can, under certain suitable conditions, be put into a superposition of quantum states that result in interference. Specifically, as we make progress in being able to observe interference phenomena with larger systems made up of multiple systems (which could form a “large” environment for each other), a conservative approach should not impose any limitations on this ability.
In a conservative approach that relies solely on decoherence, we can assume that system $S$ will only have definite values if it interacts with systems $E$ that have been decohered shortly before the interaction with $S$ begins. These systems $E$ would have been decohered by other systems that were previously decohered by some other systems, and so on. This assumption is plausible because it is reasonable to consider that the constituents of our measurement devices were previously subject to decoherence due to the large environment they interact with, and so on. However, this idea seems to start giving rise to an infinite regress and some vagueness regarding how to think about these interactions since it is unclear what decoherence even represents. Thus, even if we decide to maintain a conservative strategy, there are some issues to deal with and choices to make when interpreting QT.
@@
EnDQT will adopt the above conservative strategy but also deal with these issues by appealing to some plausible special systems that establish when these interactions began, a more precise structure that represents such interactions and simple and conservative rules that establish how determinate values arise from them. This strategy will also advance a clear view of what decoherence is about. The structure that EnDQT will introduce is a network structure whose edges represent interactions between certain systems represented via decoherence, and that roughly establishes when these interactions give rise to them having determinate values. These interactions form what I will call Stable Determination Chains (SDCs), where what I mean by “stable” will be clarified below.
Furthermore, to circumvent an infinite regress, SDCs started somewhere in spacetime. As I will argue, the first systems with determinate values arose in the past through some systems. These systems started chains of local interactions over time and space, i.e., the SDCs. More concretely, by interacting with these systems, a system $S^{\prime}$ acquires the capacity to have a determinate value of an observable during these interactions and to give rise to other systems having determinate values in interactions with $S^{\prime}$, which allows these later systems to lead other systems to have determinate values, and so on. I will call this capacity, the determination capacity. So, these chains allow the determination capacity to propagate between systems, where it is indeterministic which values will arise under interactions among the possible ones.
These interactions are modeled via decoherence; thus, as I have said, they don’t lead to any modification of fundamental equations of QT. Moreover, note that the systems that never were part of this chain of interactions can, in principle, continue their deterministic evolution represented unitarily, and without any indeterministic process interrupting this evolution. This allows EnDQT to fulfill the desideratum **).
I will argue that one possible systems that initiate SDCs can be understood through inflation, a dominant paradigm in modern cosmology. So, EnDQT will assign a new role to the systems that start inflation. I will also argue that this assumption regarding the special systems that give rise to SDCs isn’t problematic because every QT has to postulate some kind of initial conditions in the early universe to explain diverse physical phenomena (e.g., the temporal asymmetries through the so-called Past Hypothesis777[1, 22].) and the phenomena that the big bang and inflation aim to explain). However, EnDQT has the benefit of potentially grounding these special features of the early universe in fundamental aspects of QT.
Like QT, the theory of relativity has enjoyed numerous empirical successes. Thus, a conservative approach should make QT compatible with relativity and not in conflict with it. Therefore, another important desideratum for a conservative approach to QT is to make it compatible with relativity at the basic level. What I mean by at basic level is setting aside issues related with the possible need of a theory of quantum gravity (more on this below). However, as is well known and will be further explained, already at the basic level of non-relativistic QT, Bell’s theorem threatens such compatibility. Thus, another desideratum for a quantum theory is that
LC) In domains where the application of QT is established, ensure that the theory: does not conflict with relativity by favoring any specific reference frame, thus avoiding action-at-a-distance phenomena as seen in Bohmian mechanics, [^8] See, e.g., [44].; does not introduce hidden variables that result in retrocausality or superdeterminism. [^9] See, e.g., [50, 40].
@@
For EnDQT to achieve LC), first, I will argue that it can deal with Bell’s theorem by providing a local explanation of quantum correlations via Quantum Causal Models (Section 3). [^10] See, e.g., [28, 3, 3]. Second, EnDQT will adopt a perspective on quantum states where they don’t literally and directly represent some physical entity; instead, together with other elements of the theory, such as observables and networks representing SDCs, they help make inferences, gain knowledge about and indirectly represent how systems evolve and affect each other, how SDCs evolve, when systems acquire or not determinate values, how systems evolve outside interactions, etc. The fulfillment of LC) will be mainly argued on the basis of non-relativistic QT by showing that quantum correlations can be explained without being in tension with relativity. It turns out that the tension between QT and relativity already arises in the non-relativistic case. I will not focus on the locality of a theory in the sense that it respects relativistic symmetries (i.e., it allows for Lorentz covariance and, more generally, the general covariance of the physical laws). Nevertheless, I will also provide arguments for why EnDQT can be local in this sense.
@@
You may call this theory a “collapse theory,” but I resist the use of that terminology because we run the risk of running together very different interpretations of QT with very different consequences and ontological commitments. This view doesn’t reify the wavefunction. In agreement with the above desideratum, contrary to spontaneous collapse theories, which clearly reify the quantum state, there is no literal physical collapse of the wavefunction/quantum states in a superposition. Instead, there is an epistemic local state update of the original state of the target system that can be implemented upon decoherence of this system by its environmental systems that belong to SDCs under their local interactions. These interactions give rise indeterministically to these systems having a determinate value. Therefore, given that EnDQT doesn’t reify quantum states, in Bell-type scenarios, the measurements of an agent Alice on her system don’t non-locally affect the space-like separated system of agent Bob and vice versa. There isn’t such non-locality. On top of that, this view doesn’t put any limits on the superposition principle, i.e., the ability to place objects in a coherent superposition, except whatever limits decoherence models already provide us. There isn’t also a special force causing this indeterministic process to occur.
@@
Note that this view on quantum states also considers that decoherence shouldn’t be interpreted as representing a process of branching of the wave-function/quantum states but rather as a process in which, under local interactions, an environmental system that belongs to an SDC gives rise indeterministically to another system having determinate values.
A common way to fulfill UT) and LC) is by adopting a relationalist interpretation. On this view, outcomes are relative to and can vary in the same situation according to, for example, worlds, any systems, private perspectives, environments, simultaneity hyperplanes, etc. [^11] See, e.g., [95, 88, 32, 47, 33, 75]. One example of a relationalist strategy is the Many-Worlds Interpretation (MWI), in which roughly each physically possible outcome corresponds to a $world.$ However, while maintaining our conservative approach, we should aim to circumvent such a counterintuitive strategy. Thus, another desiderata is to
NR) Not adopt a relationalist interpretation of QT.
Furthermore, given the well-known issue of probabilities of the MWI, [^12] See, e.g., [2, 83]. fulfilling this desideratum presents another advantage. Also, it is unclear if relationalism in a single world is desirable. [^13] See, e.g., [86]. As it will become clear, EnDQT will fulfill this desideratum since outcomes will be absolute/non-relationalist.
Note that MWI-like views consider that decoherence in large enough regions of spacetime establishes criteria for systems to have determinate values of an observable. However, as we will see in the next section, EnDQT, in a sense, considers that such criteria form necessary but not sufficient conditions for determinate values to arise. This is because it matters if the environmental systems involved in this process belong to an SDC.
@@
Moreover, as I will argue (section 3), to my knowledge, EnDQT is currently the only QT that doesn’t modify the fundamental equations of standard QT, [^14] It might be objected that quantum causal models modify the fundamental equation of QT. Note that I don’t regard these models as a modification of equations of standard QT, but as a generalization or variation. that is able to provide a local, non-hidden variable, non-relationalist common cause explanation of quantum correlations like the ones in Bell scenarios and the so-called extended Wigner’s friend scenarios. [^15] I will focus on the scenarios from [14, 16]. [^16] Contrary to the suggestions by others (e.g., [24, 90, 106]), we don’t need to adapt QCMs to a relationalist approach, which might be considered as an advantage of this view. Thus, by fulfilling all of the above plausible desiderata, EnDQT should be considered a conservative approach, which, given the empirical success of QT, I think is a virtue.
I will start by explaining the basics of EnDQT, argue that it provides UT) and start building the case that it provides LC) and NR) (Section 2). I will also provide a toy model to illustrate how EnDQT represents physical interactions in Appendix A. In Section 3, I will argue that EnDQT provides LC) and NR) by showing that it provides a non-relational, local, and non-superdeterministic/non-retrocausal explanation of quantum/Bell-type correlations. Throughout the text, I will suggest future developments.
Therefore, I will argue that in contrast to the well-known quantum theories, EnDQT has the benefit of addressing the measurement problem without adding non-local, superdeterministic, or retrocausal hidden variables. Also, it addresses that problem without being in tension with relativistic causality by providing a local causal explanation of quantum correlations. Furthermore, measurement outcomes don’t vary according to, for example, systems or worlds. It is a conservative QT in the sense that, unlike theories such as spontaneous collapse theories, no modifications of the fundamental equations of quantum theory are required to establish when determinate values arise. Moreover, in principle, arbitrary systems can be in a superposition for an arbitrary amount of time. To simplify, I will mostly assume non-relativistic QT and the Schrödinger picture Hilbert space-based finite dimensional QT.
@@ 2 Main features: four conditions and two hypotheses
I will start by presenting the main features of EnDQT and show why it fulfills the desideratum UT), which was explained in the previous section. Also, I will begin building the argument for why it allows for LC) and NR). The main features of EnDQT presented here [^17] We will see that there are other versions of EnDQT if we vary these hypotheses and conditions, but we will see that this version so far is enough to fulfill the desiderata presented in Section 1 involve four conditions and two hypotheses about how the capacity to have determinate values, and to provide other systems that capacity, spreads through interactions. After some key assumptions, I will start by explaining the four conditions, which will require that I give an account of the role of quantum states (which was already mostly given in the previous section), systems, interactions, decoherence, and how having a determinate value and allowing other systems to have determinate values propagate via interactions inferred via decoherence, where specific chains of interactions are formed. Afterward, I will present the two hypotheses that support EnDQT, where the second hypothesis involves an account of what the systems that start SDCs could be. I will mention some empirical predictions that EnDQT provides throughout the next sections.
2.1 The four conditions